OpenStudy (anonymous):

A right triangle has one leg 24 m long and a hypotenuse 25 m long. How long is its other leg?

OpenStudy (anonymous):

leg^2 +leg^2=hypotenuse^2

OpenStudy (anonymous):

7 - 24 - 25 It's a pythagorean triple So you know the other leg is 7

OpenStudy (anonymous):

24^2 +leg^2=25^2 576+leg^2=625

OpenStudy (anonymous):

I forgot about that... but moon wolf, do u know how to solve it using pythagoreans theorem?

OpenStudy (anonymous):

a^2+b^2=c^2 ; 24^2+b^2= 25^2; 576+b^2=625; -576 -576; b^2=49; square root b^2 and 49; b=7